OpenStudy (anonymous):

solve the system using elimination -8-6x=-7y 1/2=-x+3/4y

OpenStudy (anonymous):

please I really need help

OpenStudy (anonymous):

Hi!

OpenStudy (anonymous):

Two equations are given. In method of elimination, you eliminate one of the 2 variables i.e. x and y, and the find the value of the other one. Then you substitute this value in one of the initial equations to find the other variable. We have two equations: -6x - 8 = -7y -x + 3/4y = 1/2 Because it's simpler, we will multiply all the elements in -x +3/4y = 1/2 by 6 to get the coefficient of x common in order to be eliminated. So: (-x+3/4y = 1/2)*6 -6x + 18/4y = 3 Now we have two equations: -6x - 8 = -7y -6x + 18/4y = 3 Rephrase equation 1: -6x + 7y = 8 -6x + 18/4y =3 Cancel -6x from both sides and subtract both equations, we have: 5y/2 = 5 Cross multiply, 5y = 10 y = 10/5 = 2 Put the value of y in equation 1. -6x +7(2) = 8 -6x + 14 = 8 14-8 = 6x 6 = 6x 6/6 = 1 =x So therefore, x = 1, y = 2. You can always verify your values by putting them in the equations and seeing if your left hand side answer is equal to the right. 1. -6x + 7y = 8 -6(1) + 7(2) -6 + 14 8 - CORRECT. 2. -6x + 18/4y = 3 -6(1) + 18(2)/4 -6 + 9 3 - CORRECT.

OpenStudy (anonymous):

thank you guys soooo Much

OpenStudy (anonymous):

I get it now!! You guys are awesome! Thanks for all the help!!